OpenStudy (anonymous):

What are the equations for the asymptotes of this hyperbola? x^2/9 - y^2/64 =1 a. y=3/8x, y=-3/8x b. y=8/3x, y=-8/3x c. y=√73/5x, y=-√73/5x d. y=√55/5x, y=-√55/5x

OpenStudy (anonymous):

its b option

OpenStudy (anonymous):

b. y=8/3x, y=-8/3x

OpenStudy (anonymous):

y - k = ± b x/ a